OpenStudy (anonymous):

Find a cubic function with the given zeros. 7, -3, 2

OpenStudy (tmak11):

i knew i should have taken notes that day

OpenStudy (tmak11):

one sec im going through last years math notes

OpenStudy (anonymous):

1. f(x) = x3 - 6x2 - 13x - 42 2. f(x) = x3 - 6x2 + 13x + 42 3. f(x) = x3 - 6x2 - 13x + 42 4. f(x) = x3 + 6x2 - 13 + 42 These are my options.

rishavraj (rishavraj):

if a,b and c are zeros of a cubic equation then the equation is (x - a)(x - b)(x - c)= 0

OpenStudy (tmak11):

so it would be B.?

OpenStudy (anonymous):

Thank you!!!

rishavraj (rishavraj):

u r most welcome.....:)) @KrissieG123